OpenStudy (kkutie7):

for my problem a car is stopping at the stop light I know the velocity I know how long it takes it to stop. I solved where v=0 so the stopping point and another thing along the journey considering the position vs time is a parabola. I have two times because v=0 twice. I don't understand which one to use I have a large one and a small one.

OpenStudy (ankit042):

Is there any equation?

OpenStudy (kkutie7):

\[\sf v_f=v_i+a\Delta t\] \[\sf 100=(4\Delta t^2)\Delta t -\dfrac{1}{2}4(\Delta t)^2\]

OpenStudy (kkutie7):

A driver needs to stop for a red light and slows down at a rate of 4m/s^2. it takes 100m to stop. (a) Find the time required to stop. (b) Find the initial velocity. (c) Plot x(t), v(t), and a(t). I need to be walk through this one.

OpenStudy (kkutie7):

I was solving for t and t=7.071s (which is the answer) and t=24.5s I'm just wondering why is it that one?

OpenStudy (ankit042):

I am not sure how you solved it but here is my solution v(final) = v(initial)+at a = -4 let us assume it takes ta time for car to stop v(final)=0 v(initial) = 4ta

OpenStudy (ankit042):

Now distance = v(initial)t +1/2 at^2

OpenStudy (ankit042):

if you solve you will get ta^2 =50

OpenStudy (kkutie7):

yeah that 7.071s

OpenStudy (ankit042):

100 = 4ta^2 -2ta^2

OpenStudy (ankit042):

Yes so only one positive value of t

OpenStudy (kkutie7):

humm. I didn't get the same. I got that number but I factored the t^2 out of an equation. I was left with 2t+1 too

OpenStudy (kkutie7):

Let me put all my work down so you can see and tell me where I messed up.

OpenStudy (ankit042):

Ohh i see the mistake in your equation!

OpenStudy (ankit042):

you messed up value of initial velocity

OpenStudy (kkutie7):

Oh I see it now! let me try again

OpenStudy (kkutie7):

vi=-4t 100=0+4t^2+4/2t^2 100=4t^2+2t^2 100=6t^2 I screwed up again somewhere...

OpenStudy (ankit042):

it is 1/2at^2 and a =-4 so you should get 100=0+4t^2-4/2t^2

OpenStudy (kkutie7):

100=4t^2-2t^2 100=2t^2 50=t^2 7.1=t thank you

OpenStudy (ankit042):

finally :) np!